# MMEarth: Exploring Multi-Modal Pretext Tasks For Geospatial Representation Learning

### v1

@@ 3.2 Image-level modalities
Biome. A Biome is a geographical region with specific vegetation, climate and animal life. We use the RESOLVE Ecoregions dataset [15] that consists of 14 terrestrial biomes. Ecoregion. The RESOLVE ecoregions [15] include 846 terrestrial ecoregions. These are a more fine-grained classification than biomes. For example the ecoregions Central African Mangroves and Indochina Mangroves belong to the biome Mangroves. ERA5 Temperature and Precipitation. ERA5 reanalysis data provides climate information back to 1950. Based on the Sentinel-2 observation date, we collect the corresponding mean, min, and max temperature as well as the total precipitation for the month, previous month, and year. Geolocation. We use cyclic encodings for the latitude and longitude of each image center. This ensures that for the longitude the cyclic property is maintained (transition from West to East). Date. We use a cyclic encoding of the month of the Sentinel-2 observation date as a proxy for the season.
@@ 4.1 Multi-Pretext Masked Autoencoder
The input image $x$ is divided into patches (or tokens) on which a random mask is applied to yield a masked image $x^{\prime}$. The encoder $f_{\theta}$ takes the visible patches as the input and returns an encoding $z$ for every visible patch. Here, the ConvNeXt V2 [52] architecture serves as the backbone which is implemented with sparse convolutions [9] to improve efficiency. We investigate two key modifications to adapt the ConvNeXt V2 architecture to medium-resolution satellite images. First, we adapted the patch size. The original ConvNeXt V2-MAE has been developed for pretraining on ImageNet and assumes input images of size $224\times 224$ pixels, which are split into $7\times 7$ patches of size $32\times 32$ pixels. MMEarth images are of size $128\times 128$ pixels, and we randomly crop them to $112\times 112$ pixels when feeding them into the MP-MAE encoder, which has a reduced patch size of $16\times 16$ pixels to preserve the $7\times 7$ patch layout. This adjustment is crucial as the number of patches is coupled with the optimal masking ratio. Second, we avoid early downsampling by modifying the first layer. In the original ConvNeXt V2 encoder, the first layer is a learned convolutional downsampling, which yields feature maps at 1/4 of the input resolution. While this approach might be efficient and looses only little spatial information that is not needed for ImageNet classification, it may not be optimal for medium-resolution satellite images such as Sentinel-2 and for pixel-level downstream tasks such as semantic segmentation with U-Net architectures [38], where high-resolution features can inform same-resolution representations via the skip connections. Therefore, we replace the first layer with an initial convolutional layer with kernel size 3 and stride 1 to learn feature maps at the input resolution followed by a depth-wise convolutional downsampling layer.
@@
We follow the ConvNeXt V2-MAE [52] decoder design and use shallow decoders with one block. We apply the same random mask to all pixel-level modalities and learn a separate decoder $h_{t}$ for each task. We consider $T$ tasks and treat most modalities as individual tasks, but group the climate variables and split latitude and longitude (see Fig. 2). To reconstruct the masked patches, a learnable mask token is used as a placeholder when combining the embedding tokens to get a dense 2D-input for the decoders. Therefore, the decoders are based on standard convolutions. For the image-level modalities, we introduce a global average pooling over the decoded non-visible patches before the final linear prediction layer. Hence, we follow the same strategy as for pixel-level modalities, whereas the encoder is encouraged to learn representations that provide the semantic context that is needed to decode the masked tokens.
@@ 4.2 Multi-task loss
All task targets $y_{t}$ are reconstructed by applying the same random mask to avoid that the model learns any shortcuts between the input and the targets. We follow prior work and compute the losses only over non-visible patches. We experimented with two settings for the multi-task loss function [48, 24]
@@
with task-specific loss $\mathcal{L}_{t}$ (e.g., cross entropy for classification and mean squared error for regression) and variance $\hat{\sigma}_{t}^{2}$. First, we applied equal weighting with $\hat{\sigma}_{t}^{2}=1$ for all decoders, followed by task-uncertainty weighting where $\hat{\sigma}_{t}^{2}$ is learnable for each decoder [24, 4]. The second setting was expected to aid in handling noisy targets, as task-specific weights naturally decrease for such tasks [24].
@@ 5 Experimental setup
Given a pretrained encoder, we followed standard practices and evaluated fine-tuning (FT) and linear probing (LP) performance on new downstream tasks. While both approaches learn the last linear layer of a neural network from scratch, FT updates the pretrained encoder and LP uses a frozen encoder.
@@ 5.1 Evaluation on downstream tasks
For downstream tasks, we considered all five Sentinel-2 datasets provided in the GEO-Bench [27] EO dataset collection. These datasets provide a diverse set of tasks and geographical coverage (Global, Europe, Africa). GEO-Bench provides harmonized subsets of the original datasets with comparable training set sizes and balanced class distributions. For clarity, we refer to the datasets with the number of training samples to avoid confusion with prior work using the original datasets. We summarize the characteristics of these five datasets in Table 2 and provide more details in the following sections.
@@
Table 2: Downstream task datsets. We evaluated on five different datasets which involve image-level classification (multi-label, multi-class) and pixel-level semantic segmentation. These datasets are modified versions from the GeoBench benchmark [27].
@@
Dataset Name Description Task Categories Input [px] Train/Val/Test BigEarth20k [41] Landcover classification in Europe Multi-label 43 120$\times$120 20k/1k/1k So2Sat20k [57] Global local climate zones classification Multi-class 17 32$\times$32 20k/1k/1k EuroSat16k [23] Landcover classification in Europe Multi-class 10 64$\times$64 16k/1k/1k Cashew1k [54] Cashew plantations in Benin (Africa) Segmentation 7 256$\times$256 1.3k/400/50 SAcrop3k [35] Crop-type segmentation in South Africa Segmentation 10 256$\times$256 3k/1k/1k
@@
Image classification: BigEarthNet [41] is a multi-label land cover classification problem with 43 categories and data from 10 countries in Europe. While the original dataset contains a total of 590k Sentinel-2 image samples, the GEO-Bench version, referred to as ‘BigEarth20k’, contains 20k training samples. So2Sat [57] is a multi-class land cover classification task with 17 categories that classifies local climate zones. The data is collected globally from urban areas and the original data contains 400k pairs of Sentinel-2 and Sentinel-1 image samples. The GEO-Bench version contains 20k training samples and is referred to as ‘So2Sat20k’. EuroSat [23] is a multi-class land use and land cover classification dataset covering Europe with 10 categories. The original dataset contains 27k image samples, and the GEO-Bench version ‘EuroSat16k’ has roughly 16k training samples. As evaluation metrics, we use the overall accuracy (Acc) for multi-class and the mean average precision (mAP) for multi-label classification.
@@
Semantic segmentation: Cashew Plantation [54] is semantic segmentation task with 7 categories that maps cashew plantations in Benin, Africa. The ‘Cashew1k’ consists of 1k training samples. SA-Crop-Type [35] is another semantic segmentation problem with 10 categories that maps crop type in South Africa. The ‘SAcrop3k’ consists of 3k training samples.
5.2 Implementation details
[…]
Semantic segmentation We adopt a U-Net [38] architecture to evaluate the pretrained encoders. The upsampling block consists of a nearest neighbour upsampling followed by a convolutional layer with 3$\times$3 filter kernels, layer norm and a GELU activation layer.
@@ 6 Results
6.1 How can domain specific pretraining improve representation?
Switching to domain specific data is not enough. Results from Table 3 indicate that just replacing the ImageNet pretraining data with domain specific optical satellite images does not improve downstream performance. We propose to slightly adjust the encoder design and the patch size to adapt to medium resolution satellite images like Sentinel-2 (10m pixels on the ground). Reducing the patch size from 32 to 16 pixels is crucial to adapt to the smaller image size of the medium resolution satellite images. This preserves the number of patches to be 7$\times$7 as in the original ConvNeXt V2 ImageNet setting and is related to an optimal masking ratio [53]. With these modifications, domain specific pretraining improved fine-tuning performance across all datasets compared to pretraining on ImageNet. While linear probing (LP) on BigEarth20k was also improved by 5pp, pretraining on ImageNet provides a strong LP baseline for So2Sat20k.
Multi-spectral images are beneficial. Since Sentinel-2 images provide more than RGB bands, we studied the impact of pretraining on multi-spectral images and found that this can improve downstream performance depending on the tasks. Both segmentation tasks profited most from the multi-spectral pretraining each with 3.5pp improvement over the RGB pretraining. We hypothesize that this is not due to the fact that these are segmentation tasks, but rather that ambiguous crop types can profit from the additional spectral information.
Table 3: Domain specific pretraining on optical images. Fine-tuning (FT) and linear probing (LP) performance using MAE pretrained on ImageNet (RGB), Sentinel-2 RGB (MMEarth-S2rgb), and Sentinel-2 multi-spectral (MMEarth-S2). Lowering the patch-size is crucial to adopt to the smaller medium resolution satellite images. Using all the 12 bands improves accuracy. Domain specific pretraining improves both FT and LP results. Higher is better for all metrics.
Pretrain data Image size Patch size BigEarth20k FT/LP So2Sat20k FT/LP Cashew1k FT SAcrop3k FT ImageNet 224 32 72.9/57.2 44.5/34.9 77.5 27.4 MMEarth-S2rgb 128 32 71.4/57.6 37.8/27.6 79.6 26.5 MMEarth-S2rgb 128 16 78.2/62.5 46.3/31.0 80.3 29.6 MMEarth-S2rgb 112 16 77.8/62.2 44.7/34.1 81.0 29.1 MMEarth-S2 112 16 78.6/64.5 46.2/28.6 84.5 33.3
@@ 6.2 How do multi-modal pretext tasks impact representations?
Multi-modal pretext tasks improve fine-tuning and linear probing. Using multi-pretext tasks improved both the FT and LP performance (see Table 4), except for the Cashew1k dataset. While MAE are known for good fine-tuning performance, the gap to the linear probing performance remains large [22, 52]. The representations learned from our multi-pretext objective substantially improved LP performance and provide a mechanism towards closing this performance gap. We observed an improvement of 3.7pp on BigEarth20k and 15.6pp on So2Sat20k with respect to pretraining on Sentinel-2 only. On So2Sat20k, the local climate zone classification, fine-tuning improves substantially (+9.7pp), indicating that the additional pretext tasks yield better representations that cannot be learned from MIM only. Here, we report results obtained with the task uncertainty loss. We found that both loss settings work equally well on these downstream tasks (see supplementary). The learned task uncertainties provide interesting insights, e.g. the uncertainty for the latitude (N-S) being lower than for the longitude (W-E), which could be explained by the rainfall and vegetation gradients along the North-South axis shaping the appearance of landscapes.
@@
Pixel-level and image-level pretext tasks are useful. To study the role of the two groups of pretext tasks, we ran experiments where the Sentinel-2 reconstruction task was extended with only one group (Table 4). For the image-level downstream tasks, pretraining with only image-level pretext tasks achieved comparable performance to pretraining with all multi-modal pretext tasks. For So2Sat20k we observed that the image-level tasks had a strong effect, possibly because ERA5 climate variables serve as a good proxy task, but further experiments would be required to understand the importance of individual pretext tasks. While we expected that pixel-level pretext tasks might be more beneficial for segmentation tasks this was not confirmed.
@@
Table 4: Multi-modal pretext tasks. Fine-tuning (FT) and linear probing (LP) performance for multi-pretext pretraining (MMEarth) and masked image pretraining (ImageNet, MMEarth-S2). We also present multi-pretext results with subsets of modalities using only pixel-level (MMEarth-PixelM) or only image-level (MMEarth-ImageM) modalities in addition to the Sentinel-2 optical image reconstruction. Multi-pretext pretraining improves both FT and LP results. Higher is better for all metrics.
Pretrain BigEarth20k FT/LP So2Sat20k FT/LP Cashew1k FT SAcrop3k FT ImageNet 72.9/57.2 44.5/34.9 77.5 27.4 MMEarth-S2 78.6/64.5 46.2/28.6 84.5 33.3 MMEarth-PixelM 79.3/66.4 49.5/37.0 80.4 34.1 MMEarth-ImageM 81.3/68.1 54.6/42.9 81.0 34.0 MMEarth 80.0/68.2 55.9/44.2 81.4 34.5
@@
(c) EuroSat16k
Figure 3: Label efficiency for few-shot downstream performance. Linear probing performance for varying downstream dataset sizes. MP-MAE (‘Atto’) pretrained on ImageNet, MMEarth-S2 (multi-spectral only), MMEarth (all multi-modal pretext tasks). Tiny-ImageNet is the original ConvNext V2-MAE with a ‘Tiny’ backbone pretrained on ImageNet using the pretrained weights from [52].
Multi-modal pretext tasks improve label efficiency. Real-world applications often lack large amount of labelled data. To explore the benefits of pretraining in a limited label regime, we evaluated the LP performance using stratified subsets of the fine-tuning datasets. In our most extreme few-shot setting, there were only 100 training samples (i.e. 2 samples/category in BigEarth20k). Results in Fig. 3 show that models pretrained on all modalities in MMEarth outperformed models trained on only Sentinel-2 or only ImageNet in few-shot linear probing. While pretraining on MMEarth-S2 outperforms ImageNet pretraining in fine-tuning (Table 4), this does not always hold for linear probing. While, EuroSat16k is an easy task where general image features learned on ImageNet are a strong baseline, we speculate that for So2Sat20k the texture in Sentinel-2 images alone is not a good pretraining signal.
Better parameter efficiency. Results from Fig. 3 show that our MP-MAE with an Atto encoder (3.7M parameters) pretrained with multi-modal pretext tasks outperformed a notably larger Tiny encoder (28M parameters) pretrained on ImageNet. Our encoder is not only smaller but was also only trained for 200 epochs compared to 800 epochs. This supports our motivation for creating better domain specific pretraining datasets and strategies instead of using larger models pretrained on generic image datasets like ImageNet.
Compares favourably to prior work. We compared our MP-MAE method to a small, but most closely related, subset of prior work in Table 5. While all models rely on different pretraining datasets and pretraining strategies, we show that our model compares favourably to prior work on most tasks, using fewer parameters in the encoder (3.7M for our MP-MAE versus 25M for ResNet-50).
Table 5: Comparison to prior work. We compare our multi-pretext pretraining with prior work that make use of different pretraining datasets and strategies. Our pretraining strategy achieve higher if not similar performance to other methods that use a ResNet-50 backbone. Higher is better for all metrics.
Method Pretrain data (locations) Encoder (parameters) BigEarth20k FT/LP So2Sat20k FT/LP Cashew1k FT SAcrop3k FT GASSL [3] fMoW (0.36M) ResNet-50 (25M) 70.3/64.3 44.7/38.2 79.3 21.5 SeCO [29] SeCo (0.20M) ResNet-50 (25M) 77.6/71.8 47.6/40.8 78.9 24.4 SatlasPretrain [7] SatlasPretrain (0.86M) ResNet-50 (25M) 72.3/61.1 38.5/27.5 65.5 19.6 MP-MAE (ours) MMEarth (1.24M) ConvNeXt V2-Atto (3.7M) 80.0/68.2 55.9/44.2 81.4 34.5

### v2

@@ 3.2 Image-level modalities
Biome. A biome is a geographical region with specific vegetation, climate and animal life. We use the RESOLVE Ecoregions dataset [15] that consists of 14 terrestrial biomes. Ecoregion. The RESOLVE ecoregions [15] include 846 terrestrial ecoregions. These are a more fine-grained classification than biomes. For example, the ecoregions Central African Mangroves and Indochina Mangroves belong to the biome Mangroves. ERA5 Temperature and Precipitation. ERA5 reanalysis data provides climate information back to 1950. Based on the Sentinel-2 observation date, we collect the corresponding mean, min, and max temperature as well as the total precipitation for the month, previous month, and year. Geolocation. We use cyclic encodings for the latitude and longitude of each image center. This ensures that for the longitude the cyclic property is maintained (transition from West to East). Date. We use a cyclic encoding of the month of the Sentinel-2 observation date as a proxy for the season.
@@ 4.1 Multi-Pretext Masked Autoencoder
The input image $x$ is divided into patches (or tokens) on which a random mask is applied to yield a masked image $x^{\prime}$. The encoder $f_{\theta}$ takes the visible patches as the input and returns an encoding $z$ for every visible patch. Here, the ConvNeXt V2 [53] architecture serves as the backbone which is implemented with sparse convolutions [9] to improve efficiency. We investigated two key modifications to adapt the ConvNeXt V2 architecture to medium-resolution satellite images. First, we adapted the patch size. The original ConvNeXt V2-MAE has been developed for pretraining on ImageNet and assumes input images of size $224\times 224$ pixels, which are split into $7\times 7$ patches of size $32\times 32$ pixels. MMEarth images are of size $128\times 128$ pixels, and we randomly crop them to $112\times 112$ pixels when feeding them into the MP-MAE encoder, which has a reduced patch size of $16\times 16$ pixels to preserve the $7\times 7$ patch layout. This adjustment is crucial as the number of patches is coupled with the optimal masking ratio. Second, we avoid early downsampling by modifying the first layer. In the original ConvNeXt V2 encoder, the first layer is a learned convolutional downsampling layer, which yields feature maps at 1/4 of the input resolution. While this approach might be efficient and loses only little spatial information that is not needed for ImageNet classification, it may not be optimal for medium-resolution satellite images such as Sentinel-2 and for pixel-level downstream tasks such as semantic segmentation with U-Net architectures [39], where high-resolution features can inform same-resolution representations via the skip connections. Therefore, we replace the first layer with an initial convolutional layer with kernel size 3 and stride 1 to learn feature maps at the input resolution followed by a depth-wise convolutional downsampling layer.
@@
We follow the ConvNeXt V2-MAE [53] decoder design and use shallow decoders with one block. We apply the same random mask to all pixel-level modalities and learn a separate decoder $h_{t}$ for each pretext task. We consider $T$ pretext tasks and treat most modalities as individual tasks, but group the climate variables and split latitude and longitude (see Fig. 2). To reconstruct the masked patches, a learnable mask token is used as a placeholder when combining the embedding tokens to get a dense 2D-input for the decoders. Therefore, the decoders are based on standard convolutions. For the image-level modalities, we introduce a global average pooling over the decoded non-visible patches before the final linear prediction layer. Hence, we follow the same strategy as for the pixel-level modalities, wherein the encoder is encouraged to learn representations that provide the semantic context that is needed to decode the masked tokens.
@@ 4.2 Multi-task loss
All pretext task targets $y_{t}$ are reconstructed by applying the same random mask to prevent the model from learning any shortcuts between the input and the targets. We follow prior work and compute the losses only over the non-visible patches. We experimented with two settings for the multi-task loss function [49, 24]
@@
with task-specific loss $\mathcal{L}_{t}$ (e.g., cross entropy for classification and mean squared error for regression) and variance $\hat{\sigma}_{t}^{2}$. First, we applied equal weighting with $\hat{\sigma}_{t}^{2}=1$ for all decoders, followed by task-uncertainty weighting where $\hat{\sigma}_{t}^{2}$ is learnable for each decoder [24, 4]. The second setting was expected to aid in handling noisy targets, as task-specific weights naturally decrease for such pretext tasks [24].
@@ 5 Experimental setup
In our problem setup, the downstream tasks are not known at pretraining time. Given a pretrained encoder, we follow standard practices and evaluate fine-tuning (FT) and linear probing (LP) performance on new downstream tasks. While both approaches learn the last linear layer of a neural network from scratch, FT updates the pretrained encoder and LP uses a frozen encoder. Implementation details are provided in the supplementary material.
@@ 5.1 Evaluation on downstream tasks
For downstream tasks, we considered all five Sentinel-2 datasets provided in the GEO-Bench [27] EO dataset collection [^2] We use the latest GEO-Bench version v1.0 in which datasets were class-balanced.. These datasets provide a diverse set of tasks and geographical coverage (Global, Europe, Africa). GEO-Bench provides harmonized subsets of the original datasets with comparable training set sizes and balanced class distributions. For clarity, we refer to the datasets with the number of training samples to avoid confusion with prior work using the original datasets. We summarize the characteristics of these five datasets in Table 2 and provide more details in the following sections.
@@
Table 2: Downstream task datasets. We evaluated on five different datasets which involve image-level classification (multi-label, multi-class) and pixel-level semantic segmentation. These datasets are modified versions from the GEO-Bench benchmark [27].
@@
Dataset Name Description Task Categories Input [px] Train/Val/Test BigEarth20k [42] Landcover classification in Europe Multi-label 43 120$\times$120 20k/1k/1k So2Sat20k [58] Global local climate zones classification Multi-class 17 32$\times$32 20k/1k/1k EuroSat2k [23] Landcover classification in Europe Multi-class 10 64$\times$64 2k/1k/1k Cashew1k [55] Cashew plantations in Benin (Africa) Segmentation 7 256$\times$256 1.3k/400/50 SAcrop3k [36] Crop-type segmentation in South Africa Segmentation 10 256$\times$256 3k/1k/1k
@@
Image classification: BigEarthNet [42] is a multi-label land cover classification problem with 43 categories and data from 10 countries in Europe. While the original dataset contains a total of 590k Sentinel-2 image samples, the GEO-Bench version, referred to as ‘BigEarth20k’, contains 20k training samples. So2Sat [58] is a multi-class land cover classification task with 17 categories that classifies local climate zones. The data is collected globally from urban areas and the original data contains 400k pairs of Sentinel-2 and Sentinel-1 image samples. The GEO-Bench version contains 20k training samples and is referred to as ‘So2Sat20k’. EuroSat [23] is a multi-class land use and land cover classification dataset covering Europe with 10 categories. The original dataset contains 27k image samples, and the GEO-Bench version ‘EuroSat2k’ has roughly 2k training samples.
@@
Semantic segmentation: Cashew Plantation [55] is a semantic segmentation task with 7 categories that maps cashew plantations in Benin, Africa. The ‘Cashew1k’ consists of 1k training samples. SA-Crop-Type [36] is another semantic segmentation problem with 10 categories that maps crop type in South Africa. The ‘SAcrop3k’ consists of 3k training samples.
U-Net evaluation for semantic segmentation: We adopt a U-Net [39] architecture to evaluate the pretrained encoders. These U-Nets are fine-tuned in two phases, referred to as FT in the results. First, we trained the randomly initialized U-Net decoder with the frozen pretrained encoder for 50 epochs, and then fine-tuned the full model for another 150 epochs.
Evaluation metrics: Following GEO-Bench [27], we use the micro-averaged overall accuracy (Acc.) for multi-class, the micro-averaged F1-score (F1) for multi-label classification, and the macro-averaged intersection over union (IoU) for semantic segmentation tasks.
@@ 6 Results
6.1 How can domain specific pretraining improve representations?
Switching to domain-specific data is not enough. The results from Table 3 indicate that just replacing the ImageNet pretraining data with domain-specific optical satellite images does not improve downstream performance. We propose slightly adjusting the encoder design and the patch size to adapt to medium resolution satellite images like Sentinel-2 (10m pixels on the ground). Reducing the patch size from 32 to 16 or 8 pixels, for image sizes of 112 or 56 pixels, respectively, is crucial to adapt to the smaller image size of the medium resolution satellite images. This preserves the number of patches to be 7$\times$7 as in the original ConvNeXt V2 ImageNet setting and is related to an optimal masking ratio [54]. With these modifications, domain-specific pretraining improved fine-tuning and linear probing performance across all datasets compared to pretraining on ImageNet. The improvement is smallest for FT on Cashew1k with 3pp and largest for FT on So2Sat with 12pp or 17pp, when training on image sizes of 112 and 56, respectively.
Multi-spectral images are beneficial. Since Sentinel-2 images provide 12 bands in total, we studied the impact of pretraining on multi-spectral images and found that this can improve downstream performance depending on the tasks (see Table 3). So2Sat20k is an exception where the model using only the RGB channels achieved high performance. The following multi-pretext results are based on using all Sentinel-2 bands as input.
Table 3: Domain specific pretraining on optical images. Fine-tuning (FT) and linear probing (LP) performance using MAE pretrained on ImageNet (RGB), Sentinel-2 RGB (MMEarth64-S2rgb, MMEarth-S2rgb), and Sentinel-2 multi-spectral (MMEarth64-S2). Lowering the patch size is crucial to adapt to the smaller medium resolution satellite images. Using all 12 bands improves accuracy. Domain-specific pretraining improves both FT and LP results.
[…]
@@ 6.2 How do multi-modal pretext tasks impact representations?
Multi-modal pretext tasks improve fine-tuning and linear probing. Using multi-pretext tasks improved both the FT and LP performance for all downstream tasks (see Table 4). While MAE are known for good FT performance, the gap with the LP performance remains large [22, 53]. The representations learned from our multi-pretext objective notably improved LP performance and provide a mechanism towards closing this performance gap. We observed an improvement of $\approx$5pp on BigEarth20k and 8pp on So2Sat20k with respect to pretraining on Sentinel-2 only. Here, we report results obtained with the task uncertainty loss. We found that both loss settings work equally well on these downstream tasks (see supplementary). However, the learned task uncertainties provide interesting insights, e.g. the uncertainty for the latitude (N-S) being lower than for the longitude (E-W), which could be explained by the rainfall and vegetation gradients along the North-South axis shaping the appearance of landscapes.
@@
Pixel-level and image-level pretext tasks are useful. To study the role of the two groups of pretext tasks, we ran experiments where the Sentinel-2 reconstruction task was extended with only one group (see Table 4). For LP we found that combining the tasks from both groups leads to the best performance on both datasets. For BigEarth20k using only one group did not improve LP performance over pretraining on Sentinel-2 only. For So2Sat20k image-level pretext tasks lead to slightly better LP performance, but pixel-level pretext tasks lead to better FT performance. While we had expected that pixel-level pretext tasks might be more beneficial for segmentation tasks, this was not confirmed. From these experiments we hypothesize that the semantics of the pretext tasks are more important than whether the tasks are pixel- or image-level.
Multi-modal pretext tasks improve label efficiency. Real-world applications often lack large amounts of labelled data. To explore the benefits of pretraining in a limited label regime, we evaluated the LP performance using stratified subsets of the fine-tuning datasets (GEO-Bench partitions). In the most extreme few-shot setting for BigEarth20k, there were only 200 training samples, i.e. 4–5 samples/category. For EuroSat2k the extreme few-shot setting has only 20 training samples with 2 samples/category. Results in Fig. 3 show that models pretrained with multi-modal pretext tasks on MMEarth outperformed models trained on only Sentinel-2 or only ImageNet in few-shot linear probing. While EuroSat is an easy task for supervised learning, GEO-Bench provides a rather small subset, which makes it a challenging few-shot task. For EuroSat2k, we found that general image features learned on ImageNet are a strong baseline working on par with our model trained on Sentinel-2 images only. Even in the extreme case when only using 1% of EuroSat2k, our multi-modal pretext tasks improve LP performance from 16.5% to 34% accuracy.
@@
Table 4: Multi-modal pretext tasks. Fine-tuning (FT) and linear probing (LP) performance for multi-pretext pretraining (MMEarth64, MMEarth) and masked image pretraining (ImageNet, MMEarth64-S2). We also present multi-pretext results with subsets of modalities using only pixel-level (MMEarth64-PixelM) or only image-level (MMEarth64-ImageM) modalities in addition to the Sentinel-2 optical image reconstruction. Multi-pretext pretraining improves both FT and LP results.
[…]
@@
(c) EuroSat2k
Figure 3: Label efficiency for few-shot downstream performance. Linear probing performance for varying downstream dataset sizes. MP-MAE (‘Atto’) pretrained on ImageNet, MMEarth64-S2 (multi-spectral only), MMEarth64 (all multi-modal pretext tasks).
Compares favorably to prior work. We compare our MP-MAE method to a small, but most closely related, subset of prior work in Table 5. While all models rely on different pretraining datasets and pretraining strategies, we show that our model compares favorably to prior work on most tasks. The encoder in our MP-MAE consists of 3.7M parameters, which is much smaller compared to e.g. the 25M for ResNet-50. This supports our motivation for creating better domain-specific pretraining datasets and pretraining strategies.
Table 5: Comparison to prior work. We compare our multi-pretext pretraining with prior work that makes use of different pretraining datasets and strategies. Our pretraining strategy achieves similar if not better performance compared to other methods that use a ResNet-50 backbone. We highlight the best and second best results.
[…]